The internal energy E of a system is given by E=bS^3/VN , where b is a constant and other
symbols have their usual meaning. The temperature of this system is equal to
1
Expert's answer
2020-09-23T08:59:09-0400

solution-

given data

energy of system(E)="\frac{bS^3}{VN}"


by applying the first law of thermodynamics


"TdS=dE+pdV"

and

"dE=TdS-pdV"

at the constant volume change in internal energy with respect to entropy is known as temperature

so,


"(\frac{\partial E}{\partial S})_v=T" .......eq1

by putting the value of E in equation 1


"(\frac{\partial\frac{bS^3}{VN} }{\partial S})_v=T"


and by partial derivative

temperature can be shown as below


"T=3\frac{bS^2}{VN}"